OpenStudy (anonymous):

mary leaves her house at noon , traveling in her car at 45 mi/h . Later , mary's brother joe leaves their house and travels the same direction at 60 mi/h . if joe leaves at 2:00 p.m. , at what time will he catch up with mary ?

OpenStudy (0bucchr):

simultaneous equations necessary: they will meet at a certain time. velocity = dist/time time = dist/vel when joe leave mary has gone 90m (45m/h*2h) djoe = dmary + 90 time = time dj/60 = dm/45 (dm + 90)/60 = dm/45 dm = 270 t = 270/60 = 4.5h